OpenStudy (anonymous):

Derive log(sin(x)) by rearranging. I get up to dy/dx = ln(10)sin(x)/cos(x)

OpenStudy (jtvatsim):

Could you explain how you found dy/dx = ln(10)sin(x)/cos(x). What steps did you take?

OpenStudy (anonymous):

I get up to \[\frac{dy}{dx}=\frac{\ln(10)\sin(x)}{\cos(x)}\] I don't know why I'm getting the inverse of the answer. I will type the steps

OpenStudy (jtvatsim):

OK, you are very close to the right answer! :) Let's take a look at your steps...

OpenStudy (anonymous):

y=log(sin(x)) \[10^y=sinx\] \[\frac{d *ln(10)*10^y}{dy}=\frac{d* cosx}{dx}\] \[\frac{dx}{dy}=\frac{cosx}{ln(10)*10^y}\] then more...

OpenStudy (jtvatsim):

Hold on one moment

OpenStudy (anonymous):

\[\frac{dy}{dx} =\frac{\ln(10)*\sin(x)}{\cos(x)}\]

OpenStudy (anonymous):

I'm making some stupid mistake but I don't know where

OpenStudy (jtvatsim):

I understood everything up to 10^y = sin(x) --> d*ln(10)*10^y/dy = d * cos x / dx

OpenStudy (jtvatsim):

Is that step valid?

OpenStudy (anonymous):

Oh, wow, I did a stupid :P. I've never taken the derivative of both sides to solve before. Is there a valid step that you can take to end up with dy/dx (without using the chain rule)?

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

Wait

OpenStudy (anonymous):

It should be:

OpenStudy (anonymous):

\[\ln(10)10^y*dy=\cos(x)*dx\]

OpenStudy (anonymous):

Is that right?

OpenStudy (jtvatsim):

Well, yes, but you are still using the chain rule to solve it. But it is mathematically correct. :) There is another way if you need it.

OpenStudy (jtvatsim):

You've used a sort of "sneaky" chain rule :)

OpenStudy (anonymous):

Haha, yeah, I can kind of see that :) What is this other way you speak of? :P

OpenStudy (jtvatsim):

Do you know the log conversion: \[\log_a(b) = \frac{\ln{b}}{\ln{a}}\]

OpenStudy (jtvatsim):

Have you seen/used that before?

OpenStudy (anonymous):

Yeah, just a sec, I'll try and work out how to use it myself and if not I will ask you what the method is. :) I like trying to work things out first.

OpenStudy (jtvatsim):

A very good habit. I applaud you! :D

OpenStudy (anonymous):

Does it also use the "sneaky" chain rule? I can only see a way to do it with the same steps but with 'e' as the base using that conversion, which makes the 'dy' slightly easier. Is that what you were thinking of?

OpenStudy (jtvatsim):

Ah, you're right! Let me see if I can cook up any other methods...

OpenStudy (anonymous):

Haha, okay cool. I like it! And that's all right, I've got three now. I really like the idea or rearranging and taking the derivative of both sides compared to the set chain rule, for the simpler ones at least. Thanks for your help :)

OpenStudy (jtvatsim):

No worries, you did very well... I'll have to think about this one for future problems. :)

OpenStudy (anonymous):

Haha, thanks